# Large Language Models in Mental Health Care: a Scoping Review

### v1

@@ Abstract
Objective: The growing use of large language models (LLMs) stimulates a need for a comprehensive review of their applications and outcomes in mental health care contexts. This scoping review aims to critically analyze the existing development and applications of LLMs in mental health care, highlighting their successes and identifying their challenges and limitations in these specialized fields.
Materials and Methods: A broad literature search was conducted in November 2023 using six databases (PubMed, Web of Science, Google Scholar, arXiv, medRxiv, and PsyArXiv) following the 2020 version of the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines. A total of 313 publications were initially identified, and after applying the study inclusion criteria, 34 publications were selected for the final review.
Results: We identified diverse applications of LLMs in mental health care, including diagnosis, therapy, patient engagement enhancement, etc. Key challenges include data availability and reliability, nuanced handling of mental states, and effective evaluation methods. Despite successes in accuracy and accessibility improvement, gaps in clinical applicability and ethical considerations were evident, pointing to the need for robust data, standardized evaluations, and interdisciplinary collaboration.
Conclusion: LLMs show promising potential in advancing mental health care, with applications in diagnostics, and patient support. Continued advancements depend on collaborative, multidisciplinary efforts focused on framework enhancement, rigorous dataset development, technological refinement, and ethical integration to ensure the effective and safe application of LLMs in mental health care.
@@
natural language processing, large language models, mental health, psychiatry
@@ 1 Introduction
Mental health is a paramount component of modern public health. Statistics from the US National Institute of Mental Health (NIMH) show that 22.8% of U.S. adults experienced mental illness of any type in 2021.[1, 2] Globally, mental health disorders account for 30% of the non-fatal disease burden, highlighting their prevalence as a major cause of disability as published by the World Health Organization (WHO). [3] Moreover, the WHO estimates that depression and anxiety disorders cost the global economy \$1 trillion per year in lost productivity. [4] These data emphasize the critical need for innovations in the prevention and management of mental health conditions, both to alleviate individual suffering and to reduce the substantial societal and economic burdens associated with mental health issues.
A large portion of mental health issues and their management take place within the realm of natural language.[5, 6] This encompasses a range of elements, including the evaluation of symptoms and signs associated with mental health disorders and various forms of interventions, such as talk therapies.[5, 6] In this context, the wealth of information embedded in recorded textual expressions and interactions serves as a crucial resource, enhancing research and practical applications in mental health.
Accordingly, Natural Language Processing (NLP), a branch of computer science that enables computers to utilize free text (natural language) information in meaningful ways, has shown promise as a tool to aid mental health-related tasks. This includes identifying specific mental conditions,[7] building emotion-support chatbots,[8] supporting interventions,[9] etc. These tasks utilize data from various sources, such as clinical data [10, 11] and social media data.[12, 13, 14, 15]
Large Language Models (LLMs), as the most recent advances in NLP, have further expanded its potential for innovative mental health care.[16, 17] As a type of Artificial Intelligence (AI) that understands and generates human-like fluent texts, LLMs offer multiple benefits to this field. For example, their efficient information retrieval and summarization capability from extensive datasets across diverse sources (e.g., electronic health records, mobile device interactions, social media platforms,[18] etc.) can provide insights into patient behaviors and experiences for clinicians. This may help with early intervention strategies and tailored treatment plans. In addition, through conversational interactions, they have the potential to offer a relatable and accessible means for individuals to articulate their emotional states and personal experiences.[19] This aids in both user self-expression and also potentially enhances clinical therapeutic processes if they can be successfully validated and integrated into the clinical workflows.
@@
We hope this study will effectively summarize current research, highlight key challenges and opportunities, and guide future work in applying large language models to mental health care. The rest of this section contextualizes the concepts of NLP and mental health care, providing a foundational understanding of the two fields. The rest of this section provides a quick introduction to the backgrounds of NLP and LLMs, as well as insights into LLM utilities in mental health care.
1.1 NLP and LLM
@@
Natural Language Processing (NLP) emerged in the 1950s with machine translation systems .[22] Initially, the focus was on rule-based systems that required explicit coding of language rules,[23, 24] which limited their scope and adaptability to handle the complexity and variability of natural languages. Statistical Language Models arose in the 1990s and they modeled word sequences using statistical methods.[25] These models were typically hindered by issues such as data sparsity. Neural language models, proposed in the 2000s, utilized neural network-inspired architectures (e.g., multi-layer perceptrons and recurrent neural networks) to refine the prediction of word sequences.[26] This era witnessed the emergence of distributed word representations, notably through the word2vec algorithm,[27] which shifted the focus towards more effective feature learning for varied NLP tasks. Pre-trained language models like ELMo,[28] BERT,[29] T5,[21] and GPT [30] were introduced in the late 2010s. These models leveraged advanced neural architectures, including bidirectional LSTM [31] and Transformers,[32] revolutionizing NLP with their ‘pre-training and fine-tuning’ paradigm. Starting from around 2020, NLP gradually moved toward the ’large language model (LLM)’ era. LLMs can be analogized to sophisticated, pre-built statistical models but are massive in their scales and can offer much more in terms of versatility and depth.
@@
Typical LLMs include GPT-3,[33] GPT-3.5,[34] GPT-4 [35] (ChatGPT is an application developed based on GPT-3.5 and GPT-4), LLaMA,[36] LLaMA-2 [37] and PaLM.[38] Thanks to their impressive performance in understanding and generating fluent human language, natural language "prompts" are proposed to interact with and instruct these LLMs with their internal weights (i.e. model parameters) fixed.[39] This strategy is generally referred to as "prompt-tuning," a paradigm distinct from "fine-tuning," where adjusting model weights requires training with additional data.
@@
Today, due to differing policies from developers, LLMs are divided into two categories: open-source and non-open-source. Open-source LLMs, e.g., LLaMA, can be deployed in local environments, allowing for customization with specific datasets while maintaining data privacy. This is particularly crucial in sensitive fields like mental health. In contrast, non-open-source LLMs like GPT-4 and PaLM do not share their model architectures or training details. These models are typically hosted by third-party organizations and can be accessed on their websites or through Application Programming Interfaces (APIs). While offering robust language processing capabilities, they require different data privacy and control considerations.
1.2 Potential Utilities in Mental Health Care
@@ 2.1 Search Strategy and Selection Criteria
We established specific criteria for paper inclusion: (1) The study must employ at least one LLM published after T5, as defined in,[20] and (2) it should directly address research questions of these models in mental health care settings. This model scope was chosen to ensure that our review concentrates on the latest advancements in the field, deliberately excluding early pre-trained language models such as BERT and GPT-2, which have already received extensive coverage in prior literature.
@@
We conducted comprehensive searches across several databases and registries, including ArXiv, MedRxiv, the ACM Digital Library, PubMed, Web of Science, and Google Scholar. The search strategy used the logical combination of keywords "Large Language Model" AND "mental OR psychiatry OR psychology." Whenever possible, we restricted the search scope to titles and abstracts. In databases without this functionality, we extended the search scope to the full text of the papers.
Further details regarding the search process and specific search queries can be found in Appendix 7.1. For the Google Scholar registry, we employed the same set of keywords and reviewed the first 100 studies, ordered by relevance, that met our inclusion criteria.
@@ 2.2 Screening with GPT-4
After removing duplicated articles and those without abstracts, we had 281 papers left for initial screening. Recognizing that GPT-4 has been shown to be adept in assisting with article screening and exhibits performance comparable to humans,[41, 42] we used GPT-4 as a secondary reviewer for this process. Prior to its use, we tested different prompts to optimize GPT-4’s screening capability (7.1).
@@ 2.3 Final Review
    Model Development and Fine-tuning: These studies propose new LLMs or introduce methods, such as fine-tuning and prompting, to adapt and improve existing LLMs for mental health care.
@@
    Application and Evaluation: These studies employ LLMs to evaluate their performance in mental health-related tasks in real-world applications. They may include the evaluation of LLMs on certain tasks (inference-only).
@@
Thirty-four articles met our criteria and were included in subsequent analyses. To maintain a clear focus and ensure a thorough analysis to answer the application-focused study questions, we summarized "Dataset and Benchmark" studies separately.
3 Results
Figure 2 shows the submission/publication time [^1] Papers with pre-prints have submission time entries. and type of the papers included in the final analysis. Research on LLMs in mental health care emerged in Sep 2022 and has shown a progressive increase in publication volume, with a surge in the latter half of the year, particularly noticeable in October. The majority of these papers are concentrated in the "Prompt-tuning and Application" category, which saw a rise in July. In contrast, "Model Development and Fine-tuning" studies were absent in the initial months, only to experience a marked rise in October. Notably, there have been only two papers concerning "Dataset and Benchmark", both emerging later in the year. Additionally, there is only one study [19] that addresses ethics, privacy, and other challenges, which was published mid-year.
Figure 2: Submission/Publication time and types distribution of studies concerning LLMs and mental health care.
@@ 3.1.1 Application Areas
The studies roughly cover three main areas. The first area is the deployment of conversational agents,[44, 19, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58] with 16 studies aiming to improve the models’ capacity to produce empathetic and context-aware responses. These agents are typically not tailored to specific mental disorders, instead, they cater to a broad spectrum of mental health needs. Among these studies, 12 are designed to interact directly with individuals seeking mental health support through various platforms such as personal digital companions,[19] on-demand online counseling,[45, 46, 47, 49, 50, 51, 53, 54, 56, 57, 58] and emotional support.[52] Some studies extend to specialized applications like couple therapy.[56, 57] The remaining four studies are developed to assist care providers by providing specific recommendations and analyses.[46, 48] This helps to mitigate the issue of provider shortage while enhancing the delivery of care. It’s noteworthy that some studies have taken a different approach by developing screening tools. These tools are designed to identify urgent cases [46] and distinguish between emotional experiences in individuals with different personality disorder subtypes.[55]
The second area focuses on resource enrichment. Studies in this category [59, 60, 61, 62, 63, 64] have explored LLM-generated explanations for multi-task analysis [59, 60] and the development of educational content.[63, 64] The latter includes creating hypothetical case vignettes relevant to social psychiatry [63] and personalized psychoeducation material.[64] The potential of synthetic data provided by LLMs is also being tapped into as a method for augmenting data and fine-tuning downstream, particularly in augmenting clinical questionnaires to enrich the symptomatology of depression cases.[62]
The third area involves the use of LLMs as classification models for detailed diagnosis.[65, 66, 59, 60, 67, 68, 69] This often involves binary classification to detect the presence or absence of a single condition providing given contexts,[66, 60, 68, 69] and multi-class classification for more granular information about a condition, including severity levels and subtypes.[65, 66, 59, 60, 67] Section 3.1.2 provides more information about the mental health conditions explored in these studies. Some examples of multi-class classification include depression severity prediction (minimal, mild, moderate, and severe according to DSM-5 [70]), subtypes of suicide (supportive, indicator, ideation, behavior, and attempt according to the Columbia Suicide Severity Rating Scale (C-SSRS) [71]), and identification of sources of stress span (school, finance, family, social relation, etc., based on the classifications established in SAD.[72]
@@ 3.1.2 Mental Health Problems Involved
A wide range of mental health conditions has been studied, but most of them are limited by the availability of open-source datasets. The most frequently studied conditions include stress,[65, 66, 59, 73, 60, 46, 47, 52, 57] suicide,[65, 66, 59, 60, 48, 68, 69] and depression.[65, 66, 59, 60, 74, 49, 75, 62, 76, 77] Other conditions include anxiety,[60, 49] bipolar disorder,[60] Post-traumatic demoralization syndrome (PTDS),[60, 78] Autism spectrum disorder (ASD),[53] and personality disorder,[55] loneliness,[60, 68] cognitive distortion,[61] In addition to these conditions, life circumstances that can profoundly affect mental health are also studied, covering relationship problems,[68, 56, 57] sleep problems, sexual violence, custody issues, and bullying.[68]
Table 1: Summary of existing LLMs for mental health, in terms of their model development, the number of parameters, the availability of training data, the training strategy, and the available public links. "B" denotes billion. TFP and IFT stand for "tuning-free prompting" and "instruction fine-tuning", respectively.
Models Base LLM # Params Base Model Training Data
@@
ChatCounselor [47] LLaMA[36] 7B Partially known IFT
@@
ExTES-LLaMA [52] LLaMA 7B Public IFT
Explainable Depression Detection [75] Vicuna[80]/GPT-3.5 13B/Unknown Unknown TFP
@@
Female Firearm Suicides [68] FLAN-UL2[82] 20B Partially known TFP
Plasticity [55] GPT-3.5 Unknown Unknown TFP
Depression Detection [76] GPT-3 Unknown Unknown TFP
Psychiatric [77] GPT-4 Unknown Unknown TFP
@@
Social psychiatry [63] GPT-3.5 Unknown Unknown TFP
Relationship counseling [56] GPT-3.5 Unknown Unknown TFP
Relationship expert [57] GPT-4 Unknown Unknown TFP
Mindfulness [64] GPT-3 Unknown Unknown TFP
Mental Wellbeing [58] GPT-3 Unknown Unknown TFP

### v3

@@ Abstract
Objectieve: This review aims to deliver a comprehensive analysis of Large Language Models (LLMs) utilization in mental health care, evaluating their effectiveness, identifying challenges, and exploring their potential for future application.
Materials and Methods: A systematic search was performed across multiple databases including PubMed, Web of Science, Google Scholar, arXiv, medRxiv, and PsyArXiv in November 2023. The review includes all types of original research, regardless of peer-review status, published or disseminated between October 1, 2019, and December 2, 2023. Studies were included without language restrictions if they employed LLMs developed after T5 and directly investigated research questions within mental health care settings.
Results: Out of an initial 313 articles, 34 were selected based on their relevance to LLMs applications in mental health care and the rigor of their reported outcomes. The review identified various LLMs applications in mental health care, including diagnostics, therapy, and enhancing patient engagement. Key challenges highlighted were related to data availability and reliability, the nuanced handling of mental states, and effective evaluation methods. While LLMs showed promise in improving accuracy and accessibility, significant gaps in clinical applicability and ethical considerations were noted.
Conclusion: LLMs hold substantial promise for enhancing mental health care. For their full potential to be realized, emphasis must be placed on developing robust datasets, development and evaluation frameworks, ethical guidelines, and interdisciplinary collaborations to address current limitations.
@@
Artificial Intelligence, Mental Health, Digital Psychiatry, Natural Language Processing
@@ 1 Introduction
Mental health is a paramount component of modern public health. Statistics from the US National Institute of Mental Health (NIMH) show that 22.8% of U.S. adults experienced mental illness of any type in 2021 [1]. Globally, mental health disorders account for 30% of the non-fatal disease burden, highlighting their prevalence as a major cause of disability as published by the World Health Organization (WHO) [2]. Moreover, the WHO estimates that depression and anxiety disorders cost the global economy \$1 trillion per year in lost productivity [3]. These data emphasize the critical need for innovations in the prevention and management of mental health conditions, both to alleviate individual suffering and to reduce the substantial societal and economic burdens associated with mental health issues.
A large portion of mental health issues and their management take place within the realm of natural language. This encompasses a range of elements, including the evaluation of symptoms and signs associated with mental health disorders and various forms of interventions, such as talk therapies. In this context, the wealth of information embedded in recorded textual expressions and interactions serves as a crucial resource, enhancing research and practical applications in mental health.
Accordingly, Natural Language Processing (NLP), a branch of computer science that enables computers to process unstructured text (natural language) information in meaningful ways, has shown promise as a tool to aid mental health-related tasks. This includes identifying specific mental conditions[4], supporting interventions [5], and building emotion-support chatbots [6]. These tasks utilize data from various sources, such as clinical data [7, 8] and social media data[9, 10, 11, 12].
Large Language Models (LLMs), one of the most recent advances in NLP, have further expanded the potential for innovative mental health care[13, 14]. As a type of Artificial Intelligence (AI) that understands and generates human-like fluent texts, LLMs offer many promising applications for mental health. For example, their ability to efficiently retrieve and summarize information from large volumes of unstructured text contained in diverse sources (e.g., electronic health records, mobile device interactions, social media platforms, etc.) can provide insights into patient behaviors and experiences for clinicians. This may help with early intervention strategies and tailored treatment plans. In addition, through conversational interactions, they have the potential to offer a relatable and accessible means for individuals to articulate their emotional states and personal experiences[15]. This aids in both user self-expression and also potentially enhances clinical therapeutic processes if they can be successfully validated and integrated into the clinical workflows.
@@
1.1 Background on Natural Language Processing and Large Language Models
@@
Natural Language Processing (NLP) emerged in the late 1940s with machine translation systems [18]. Initially, the focus was on rule-based systems that required explicit coding of language rules[19, 20], which limited their scope and adaptability to handle the complexity and variability of natural languages. Statistical Language Models arose in the 1990s and they modeled word sequences using statistical methods[21]. These models were typically hindered by issues such as data sparsity. Neural language models, proposed in the 2000s, utilized neural network-inspired architectures (e.g., multi-layer perceptrons and recurrent neural networks) to refine the prediction of word sequences[22]. This era witnessed the emergence of distributed word representations, notably through the word2vec algorithm[23]. which shifted the focus towards more effective feature learning for varied NLP tasks. Pre-trained language models like ELMo[24], BERT[25], T5[17], and GPT [26] were introduced in the late 2010s. These models leveraged advanced neural architectures, including bidirectional LSTM [27] and Transformers[28], revolutionizing NLP with their ‘pre-training and fine-tuning’ paradigm. Starting from around 2020, NLP gradually moved toward the ’large language model (LLM)’ era. LLMs can be analogized to sophisticated, pre-built statistical models but are massive in their scales and can offer much more in terms of versatility and depth.
@@
Typical LLMs include GPT-3[29], GPT-3.5[30], GPT-4 [31] (ChatGPT is an application developed based on GPT-3.5 and GPT-4), LLaMA[32], LLaMA-2 [33] and PaLM[34]. Thanks to their impressive performance in understanding and generating fluent human language, natural language ”prompts” are proposed to interact with and instruct these LLMs with their internal weights (i.e. model parameters) fixed[35]. This strategy is generally referred to as ”prompt-tuning,” a paradigm distinct from ”fine-tuning,” where adjusting model weights requires training with additional data.
@@
Today, due to differing policies from developers, LLMs are divided into two categories: open-source and non-opensource. Open-source LLMs, e.g., LLaMA, can be deployed in local environments, allowing for customization with specific datasets. This local deployment ensures data privacy is maintained, as it obviates the necessity for data to be uploaded to cloud-based services, which is crucial in sensitive fields like mental health. In contrast, non-open-source LLMs like GPT-4 and PaLM do not share their model architectures or training details. These models are typically hosted by third-party organizations and can be accessed on their websites or through Application Programming Interfaces (APIs). While offering robust language processing capabilities, they require different data privacy and control considerations.
1.2 Potential Applications in Mental Health Care
@@ 2.1 Search Strategy and Selection Criteria
We established specific criteria for paper inclusion: (1) The study must employ at least one LLM published after T5, as defined in[16], and (2) it should directly address research questions of these models in mental health care settings. This model scope was chosen to ensure that our review concentrates on the latest advancements in the field, deliberately excluding early pre-trained language models such as BERT and GPT-2, which have already received extensive coverage in prior literature.
@@
We conducted comprehensive searches across several databases and registries, including ArXiv, MedRxiv, the ACM Digital Library, PubMed, Web of Science, and Google Scholar. The search strategy used the logical combination of keywords ”Large Language Model” AND ”mental OR psychiatry OR psychology.” Whenever possible, we restricted the search scope to titles and abstracts. In databases without this functionality, we extended the search scope to the full text of the papers.
Further details regarding the search process and specific search queries can be found in Appendix B.1. For the Google Scholar registry, we employed the same set of keywords and reviewed the first 100 studies, ordered by relevance, that met our inclusion criteria.
@@ 2.2 Screening with GPT-4
After removing duplicated articles and those without abstracts, we had 281 papers left for initial screening. Recognizing that GPT-4 has been shown to be adept in assisting with article screening and exhibits performance comparable to humans[37, 38], we used GPT-4 as a secondary reviewer for this process. Prior to its use, we tested different prompts to optimize GPT-4’s screening capability (B.1).
@@ 3 Findings
    Methods Development: These studies propose new LLMs or introduce methods, such as fine-tuning and prompting, to adapt and improve existing LLMs for mental health care.
@@
    Specific Applications: These studies employ LLMs to evaluate their performance in mental health-related tasks in real-world applications. They may include the evaluation of LLMs on certain tasks (inference-only).
@@
Thirty-four articles met our criteria and were included in subsequent analyses. To maintain a clear focus and ensure a thorough analysis to answer the application-focused study questions, we summarized ”Dataset and Benchmark” studies separately.
Figure 2 shows the submission/publication time [^1] Papers with pre-prints have submission time entries. and type of the papers included in the final analysis. Research on LLMs in mental health care emerged in Sep 2022 and has shown a progressive increase in publication volume, with a surge in the latter half of the year, particularly noticeable in October. The majority of these papers are concentrated in the ”Prompt-tuning and Application” category, which saw a rise in July. In contrast, ”Model Development and Fine-tuning” studies were absent in the initial months, only to experience a marked rise in October. Notably, there have been only two papers concerning ”Dataset and Benchmark”, both emerging later in the year. Additionally, there is only one study [15] that addresses ethics, privacy, and other challenges, which was published mid-year.
Figure 2: Submission/Publication time and type distribution of qualified studies.
@@ 3.1.1 Application Areas
The studies roughly cover three main areas. The first area is the deployment of conversational agents (also called ”chatbots”) [40, 15, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54], with 16 studies aiming to improve the models’ capacity to produce empathetic and context-aware responses. These agents are typically not tailored to specific mental disorders, instead, they cater to a broad spectrum of mental health needs. Among these studies, 12 are designed to interact directly with individuals seeking mental health support through various platforms such as personal digital companions [15], on-demand online counseling [41, 42, 43, 45, 46, 47, 49, 50, 52, 53, 54], and emotional support [48]. Some studies extend to specialized applications like couple therapy [52, 53]. The remaining four studies are developed to assist care providers by providing specific recommendations and analyses[42, 44]. This helps to mitigate the issue of provider shortage while enhancing the delivery of care. It’s noteworthy that some studies have taken a different approach by developing screening tools. These tools are designed to identify urgent cases [42] and distinguish between emotional experiences in individuals with different personality disorder subtypes[51].
The second area focuses on resource enrichment. Studies in this category [55, 56, 57, 58, 59] have explored LLM-generated explanations for multi-task analysis [55, 56] and the development of educational content[58, 59]. The latter includes creating hypothetical case vignettes relevant to social psychiatry [58] and personalized psychoeducation material[59]. The potential of synthetic data provided by LLMs is also being tapped into as a method for augmenting data and fine-tuning downstream, particularly in augmenting clinical questionnaires to enrich the symptomatology of depression cases[57].
The third area involves the use of LLMs as classification models for detailed diagnosis[60, 61, 55, 56, 62, 63, 64]. This often involves binary classification to detect the presence or absence of a single condition providing given contexts[61, 56, 63, 64], and multi-class classification for more granular information about a condition, including severity levels and subtypes[60, 61, 55, 56, 62]. Section 3.1.2 provides more information about the mental health conditions explored in these studies. Some examples of multi-class classification include depression severity prediction (minimal, mild, moderate, and severe according to DSM-5 [65]), subtypes of suicide (supportive, indicator, ideation, behavior, and attempt according to the Columbia Suicide Severity Rating Scale (C-SSRS) [66]), and identification of sources of stress span (school, finance, family, social relation, etc., based on the classifications established in SAD[67].
@@ 3.1.2 Mental Health Problems Involved
A wide range of mental health conditions has been studied, but most of them are limited by the availability of opensource datasets. The most frequently studied conditions include stress[60, 61, 55, 68, 56, 42, 43, 48, 53], suicide[60, 61, 55, 56, 44, 63, 64], and depression[60, 61, 55, 56, 69, 45, 70, 57, 71, 72]. Other conditions include anxiety[56, 45], bipolar disorder[56], Post-traumatic demoralization syndrome (PTDS)[56, 73], Autism spectrum disorder (ASD)[49], and personality disorder[51], loneliness[56, 63], cognitive distortion[74], In addition to these conditions, life circumstances that can profoundly affect mental health are also studied, covering relationship problems[63, 52, 53], sleep problems, sexual violence, custody issues, and bullying[63].
Table 1: Summary of existing LLMs for mental health, in terms of their model development, the number of parameters, the availability of training data, the training strategy, and the available public links. ”B” denotes billion. TFP and IFT stand for ”tuning-free prompting” and ”instruction fine-tuning”, respectively. A Base Model is a foundational model trained on a general dataset, which serves as the starting point for fine-tuning. Therefore, in the ”Base Model” column, a model like ChatGPT would list its foundation, which is GPT-3.5.
Method Base LLM # Params Base Model Training Data
@@
ChatCounselor [43] Vicuna[77] 7B Partially knwon IFT
@@
Zheng et al. [48] LLaMA[32] 7B Public IFT
Bao et al. [70] Vicuna/GPT-3.5/LLaMA/Mistral[78] 13B/Unknown/33B/7B Unknown TFP/IFT
@@
Zhou et al. [63] FLAN-UL2[80] 20B Partially known TFP
Hadar Shoval et al. [51] GPT-3.5 Unknown Unknown TFP
Hayati et al. [71] GPT-3 Unknown Unknown TFP
Grabb [72] GPT-4 Unknown Unknown TFP
@@
Smith et al. [58] GPT-3.5 Unknown Unknown TFP
Vowels et al. [52] GPT-3.5 Unknown Unknown TFP
Vowels [53] GPT-4 Unknown Unknown TFP
Kumar et al. [59] GPT-3 Unknown Unknown TFP
Kumar et al. [54] GPT-3 Unknown Unknown TFP